# Center-freeness of finite-step solvable groups arising from anabelian geometry

### v1

@@ 2.1. A strategy for $m$-step solvable center-freeness
The center-freeness of $G$ follows from ab-faithfulness. The center-freeness of $G^{m}$ follows from Lemmas 2.3 and 2.4. ∎
$2.$ Proof of the center-freeness of the maximal $m$-step solvable quotients of the geometric étale fundamental groups of hyperbolic curves
2.2.1.
For any connected, locally Noetherian scheme $\mathcal{X}$, we define
$\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X},\ast)$
to be the étale fundamental group of $\mathcal{X}$, where $\ast:\operatorname{Spec}(\Omega)\rightarrow\mathcal{X}$ denotes a geometric point of $\mathcal{X}$ and $\Omega$ denotes an algebraically closed field. The profinite group $\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X},\ast)$ depends on the choice of base point only up to inner automorphisms, and therefore we omit the choice of base point below. If, moreover, $\mathcal{X}$ is a geometrically connected $k$-scheme for some field $k$ of characteristic 0, then we set
$\Delta_{\mathcal{X}}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X}_{\overline{k}})^{\Sigma},\qquad\text{and}\qquad\Pi_{\mathcal{X}}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X},\ast)/\ker(\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X}_{\overline{k}})\to\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X}_{\overline{k}})^{\Sigma}).$
In this case, we have the following exact sequence, called the homotopy exact sequence:
$1\to\Delta_{\mathcal{X}}\to\Pi_{\mathcal{X}}\to G_{k}\to 1.$
Additionally, we define
$\Pi_{\mathcal{X}}^{(m)}\coloneq\Pi_{\mathcal{X}}/\Delta_{\mathcal{X}}^{[m]}.$
Then the homotopy exact sequence naturally induces the following exact sequence:
$1\to\Delta_{\mathcal{X}}^{m}\to\Pi_{\mathcal{X}}^{(m)}\to G_{k}\to 1.$ (2.4)
2.2.2.
Next, we consider the case of smooth curves, where we always assume that smooth curves are geometrically connected. We say that a smooth curve $X$ over $k$ of type $(g,r)$ is hyperbolic if 22gr<02-2g-r<0, i.e., $(g,r)\notin\{(0,0),(0,1),(0,2),(1,0)\}$. When $r\geq 1$, we say that $X$ is affine. The basic fact about hyperbolicity is that
[…]
(see [14, Corollary 1.4]). If $X$ is affine, then $\Delta_{X}$ is a free pro-$\Sigma$ group; if $X$ is proper, then $\Delta_{X}$ is a pro-$\Sigma$ surface group.
@@
Let $X$ be a hyperbolic curve over a field $k$ of characteristic 0. Let $m\in\mathbb{Z}_{\geq 2}$. Then $\Delta_{X}^{m}$ is center-free. If, moreover, $G_{k}$ is center-free, then $\Pi_{X}^{(m)}$ is also center-free.
@@
The last assertion follows from the first assertion and 2.4. Then we show the first assertion. We may assume $k=\overline{k}$. Since $X$ is hyperbolic, $\Delta_{X}$ is non-abelian. When $X$ is affine, $\Delta_{X}$ is a free pro-$\Sigma$ group and hence the assertion follows from Corollary 1.6. Therefore, we may assume that $X$ is proper, i.e., $\Delta_{X}$ is a pro-$\Sigma$ surface group of genus $g\geq 2$. By Proposition 2.5, it suffices to show that $\Delta_{X}$ is ab-torsion-free and ab-faithful. The known result [14, Corollary 1.2] implies that $\Delta_{X}^{\mathrm{ab}}$ is torsion-free and hence $\Delta_{X}$ is ab-torsion-free, since any open subgroup of $\Delta_{X}$ is also an étale fundamental group of a proper hyperbolic curve over $k$.
Next, we show that $\Delta_{X}$ is ab-faithful. Let $H$ be an open subgroup of $\Delta_{X}$ and NN an open normal subgroup of $H$. To prove ab-faithfulness, we may replace $\Delta_{X}$ by $H$ and assume that $H=\Delta_{X}$. Let $X_{N}\to X$ be the connected finite étale Galois covering corresponding to NN, with Galois group $\operatorname{Gal}(X_{N}/X)(\cong\Delta_{X}/N)$. Fix $\ell\in\Sigma$. Then the morphism
$\operatorname{Gal}(X_{N}/X)\rightarrow\operatorname{Aut}_{\mathbb{Q}_{\ell}}(H^{1}_{\mathrm{\acute{e}t}}(X_{N},\mathbb{Q}_{\ell}))$
is also injective. (This is a classical result; see, for instance, [7, Proposition 1.3].) On the other hand, the $\operatorname{Gal}(X_{N}/X)$-module $H^{1}_{\mathrm{\acute{e}t}}(X_{N},\mathbb{Q}_{\ell})$ is the $\mathbb{Q}_{\ell}$-linear dual of $N^{\mathrm{ab},\ell}\otimes_{\mathbb{Z}_{\ell}}\mathbb{Q}_{\ell}$, with the conjugation action of $\Delta_{X}/N$ (see [4, Exposé XI, Section 5]). Therefore, the composition of the natural morphisms
$\Delta_{X}/N\to\operatorname{Aut}(N^{\mathrm{ab}})\to\operatorname{Aut}_{\mathbb{Z}_{\ell}}(N^{\mathrm{ab},\ell})\to\operatorname{Aut}_{\mathbb{Q}_{\ell}}\bigl(N^{\mathrm{ab},\ell}\otimes_{\mathbb{Z}_{\ell}}\mathbb{Q}_{\ell}\bigr).$
Remark 2.7.
Let $k$ be a field finitely generated over $\mathbb{Q}$. Then $G_{k}$ is center-free [3, Proposition 19.2.6], and hence $\Pi_{X}^{(m)}$ is center-free for all $m\in\mathbb{Z}_{\geq 2}$ by Theorem 2.6. On the other hand, $\Pi_{X}^{(\mathrm{ab})}$ is also center-free, although $\Delta_{X}^{\mathrm{ab}}$ is not; see [16, Proposition 1.3] for details.
@@
The maximal $m$-step solvable quotients of a pro-$\Sigma$ surface group of genus $g\geq 2$ are center-free for all $m\in\mathbb{Z}_{\geq 2}$.
@@ 2.3. Injectivity of the $m$-step solvable Grothendieck conjecture
Let $i$ range over $\{1,2\}$. Let $m\in\mathbb{Z}_{\geq 1}$. Let $k$ be a field of characteristic 0 with algebraic closure $\overline{k}$, and let $X_{i}$ be a smooth curve over $k$. We write $\tilde{X}_{i}^{m}\to X_{i}$ for the maximal geometrically $m$-step solvable pro-$\Sigma$ Galois covering of $X_{i}$, which is a scheme over $\overline{k}$. We introduce the following non-standard notation for isomorphism sets:
@@
    $\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)$
@@
    $\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m+n)},\Pi_{X_{2}}^{(m+n)}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr).$
    Moreover, we define
    $\operatorname{Isom}_{G_{k}}^{\operatorname{Out},(m+n)}(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}})\coloneq\operatorname{Isom}_{G_{k}}^{(m+n)}(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}})/\operatorname{Inn}(\Delta_{X_{2}}^{m}),$
    where $\operatorname{Inn}(\Delta_{X_{2}}^{m})$ denotes the subgroup of $\operatorname{Isom}_{G_{k}}(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}})$ consisting of inner automorphisms induced by conjugation by elements of $\Delta_{X_{2}}^{m}$.
@@
Theorem ([8, Theorem 18.1]).
Assume $\Sigma=\{p\}$. Let $i$ range over $\{1,2\}$. Let $m\in\mathbb{Z}_{\geq 2}$. Let $k$ be a sub-$p$-adic field (i.e., a field that embeds as a subfield of a finitely generated extension of $\mathbb{Q}_{p}$) with algebraic closure $\overline{k}$, and let $X_{i}$ be a smooth curve over $k$. Assume that at least one of $X_{1}$ and $X_{2}$ is hyperbolic. Then the natural map
$\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)$ (2.5)
@@
The following theorem shows that the natural map is also injective:
@@
We keep the notation and assumptions as in the above theorem. Then the natural map 2.5 is bijective.
@@
If $\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)$$\emptyset$, then the statement is tautological. Hence we may assume that $\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)\neq\emptyset$. First, by Theorem 2.6, $\Delta_{X_{1}}^{m}$ is nontrivial and center-free if $X_{1}$ is hyperbolic. If $X_{1}$ is not hyperbolic, then $\Delta_{X_{1}}^{m}$ is abelian. Therefore, we can reconstruct whether $X_{1}$ is hyperbolic from $\Delta_{X_{1}}^{m}$. Hence we may assume that $X_{1}$ and $X_{2}$ are both hyperbolic. Next, by definition, there is an exact sequence:
@@
$1\to\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\to\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\to\operatorname{Isom}_{k}(X_{1},X_{2})\to 1.$
@@
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{k}(X_{1},X_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Inn}(\Delta_{X_{2}}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{G_{k}}^{\operatorname{Out},(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
By the definition of $\tilde{X}_{2}^{m}\to X_{2}$, we have a canonical identification
$\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\,\cong\,\Delta_{X_{2}}^{m}.$
By Theorem 2.6, $\Delta_{X_{2}}^{m}$ is center-free. Therefore,
$\ker(\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\twoheadrightarrow\operatorname{Inn}(\Delta_{X_{2}}^{m}))=\mathrm{C}_{\Pi_{X_{2}}^{(m)}}(\Delta_{X_{2}}^{m})$
is trivial. Hence the left-hand vertical arrow in the above commutative diagram is bijective. Moreover, the right-hand vertical arrow is surjective by [8, Theorem 18.1], and injective by [16, Lemma 4.9]. (Note that [16, Lemma 4.9] assumed that $k$ is a field finitely generated over $\mathbb{Q}$. However, the proof can be applied to the case where $k$ is a sub-$p$-adic field.) Thus, by the snake lemma, the middle vertical arrow is also bijective. This completes the proof. ∎
@@
In Theorem 2.9, we assumed that $\Sigma=\{p\}$, i.e., $\Sigma=1$. This assumption is not needed if we further assume $m\geq 3$. The author expects that, even for $m=2$, the same statement should hold for an arbitrary $\Sigma$. To prove this, we would need to check whether the proof of [8, Theorem 18.1] applies in this setting as well. At the time of writing, the author has not attempted this modification.

### v3

@@ 2.1. Ab-torsion-freeness and ab-faithfulness
For simplicity, we write
$\Delta_{X}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(X)^{\Sigma}\qquad(\text{resp. }\pi_{1}^{\mathrm{tame}}(X)^{\Sigma}).$
The known result [14, Corollary 1.2] implies that any open subgroup of $\Delta_{X}$ is torsion-free, since any open subgroup is also an étale fundamental group of a hyperbolic curve over $k$. Hence $\Delta_{X}$ is ab-torsion-free.
Next, we show the ab-faithfulness. Let $H$ be an open subgroup of $\Delta_{X}$ and $N$ an open normal subgroup of $H$. To prove ab-faithfulness, we may replace $\Delta_{X}$ by $H$ and assume that $H=\Delta_{X}$. Let $Y\to X$ be the connected finite étale Galois covering corresponding to $N$, with Galois group $\Gamma\coloneq\Delta_{X}/N$. Let $\ell\in\Sigma$ be a prime number different from the characteristic of $k$. Then the action
$\Gamma\curvearrowright\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})$
is faithful by Lemma 2.7. On the other hand, the $\Gamma$-module $\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})$ is the $\mathbb{Q}_{\ell}$-linear dual of $N^{\mathrm{ab},\ell}\otimes_{\mathbb{Z}_{\ell}}\mathbb{Q}_{\ell}$, with the conjugation action of $\Gamma$ (see [3, Exposé XI, Section 5]). Therefore, the composition of the natural morphisms
$\Gamma=\Delta_{X}/N\to\operatorname{Aut}(N^{\mathrm{ab}})\to\operatorname{Aut}_{\mathbb{Z}_{\ell}}(N^{\mathrm{ab},\ell})\to\operatorname{Aut}_{\mathbb{Q}_{\ell}}\bigl(N^{\mathrm{ab},\ell}\otimes_{\mathbb{Z}_{\ell}}\mathbb{Q}_{\ell}\bigr).$
2.2.3.
The following is the first main theorem of this paper:
@@
Let $X$ be a hyperbolic curve over a field $k$. Assume that $k$ is a separably closed field and that $\Sigma$ contains a prime number different from the characteristic of $k$. Then, for any $m\in\mathbb{Z}_{\geq 2}$, the maximal $m$-step solvable quotients of $\pi_{1}^{\mathrm{\acute{e}t}}(X)^{\Sigma}$ and $\pi_{1}^{\mathrm{tame}}(X)^{\Sigma}$ are both torsion-free and center-free.
@@
The assertion follows from Propositions 2.5 and 2.8. ∎
@@
For any $m\in\mathbb{Z}_{\geq 2}$, the maximal $m$-step solvable quotients of a pro-$\Sigma$ surface group of genus $g\geq 2$ are torsion-free and center-free.
@@ 2.3. Injectivity of the $m$-step solvable Grothendieck conjecture
Next, we explain an application of Theorem 2.9 to the $m$-step solvable analogue of the Grothendieck conjecture. Let $X$ be a smooth curve over a field $k$. In this subsection, we focus only on the case where the field $k$ has characteristic 0 (or, more restrictively, the field $k$ is a sub-$p$-adic field for some prime number $p$, i.e., a field that embeds as a subfield of a finitely generated extension of $\mathbb{Q}_{p}$). For simplicity, we set
$\Delta_{X}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{k}})^{\Sigma},\qquad\text{and}\qquad\Pi_{X}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(X)/\ker(\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{k}})\to\Delta_{X}),$
where $\overline{k}$ is an algebraic closure of $k$. In this notation, we have the following exact sequence, called the homotopy exact sequence:
$1\to\Delta_{X}\to\Pi_{X}\to G_{k}\to 1.$
We also define
$\Pi_{X}^{(m)}\coloneq\Pi_{X}/\Delta_{X}^{[m]}.$
By construction, the homotopy exact sequence naturally induces the following exact sequence:
$1\to\Delta_{X}^{m}\to\Pi_{X}^{(m)}\to G_{k}\to 1.$ (2.4)
2.3.2.
Let $i$ range over $\{1,2\}$. Let $m\in\mathbb{Z}_{\geq 1}$. Let $X_{i}$ be a smooth curve over $k$. We write $\widetilde{X_{i}}^{m}\to X_{i}$ for the maximal geometrically $m$-step solvable pro-$\Sigma$ Galois covering of $X_{i}$, which is a scheme over $\overline{k}$. For this, we introduce the following non-standard notation for isomorphism sets:
@@
    $\operatorname{Isom}_{\overline{k}/k}\bigl(\widetilde{X_{1}}^{m}/X_{1},\widetilde{X_{2}}^{m}/X_{2}\bigr)$
@@
    $\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m+n)},\Pi_{X_{2}}^{(m+n)}\bigr)\to\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr).$
    We also define
    $\operatorname{Isom}_{G_{k}}^{\operatorname{Out},(m+n)}(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)})\coloneq\operatorname{Isom}_{G_{k}}^{(m+n)}(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)})/\operatorname{Inn}(\Delta_{X_{2}}^{m}),$
    where $\operatorname{Inn}(\Delta_{X_{2}}^{m})$ denotes the subgroup of $\operatorname{Isom}_{G_{k}}(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)})$ consisting of inner automorphisms induced by conjugation by elements of $\Delta_{X_{2}}^{m}$.
@@
Theorem ([7, Theorem 18.1]).
Assume $\Sigma=\{p\}$. Let $i$ range over $\{1,2\}$. Let $m\in\mathbb{Z}_{\geq 2}$. Let $k$ be a sub-$p$-adic field with algebraic closure $\overline{k}$, and let $X_{i}$ be a smooth curve over $k$. Assume that at least one of $X_{1}$ and $X_{2}$ is hyperbolic. Then the natural map
$\operatorname{Isom}_{\overline{k}/k}\bigl(\widetilde{X_{1}}^{m}/X_{1},\widetilde{X_{2}}^{m}/X_{2}\bigr)\to\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)$ (2.5)
@@
2.3.3.
The following is the second main theorem of this paper:
@@
We keep the notation and assumptions as in the above theorem. Then the natural map (2.5) is bijective.
@@
If $\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)=\emptyset$, then the statement is tautological. Hence we may assume that $\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\neq\emptyset$. First, by Theorem 2.9, the group $\Delta_{X_{1}}^{m}$ is nontrivial and center-free if $X_{1}$ is hyperbolic. If $X_{1}$ is not hyperbolic, then $\Delta_{X_{1}}^{m}$ is abelian. Therefore, we can determine whether $X_{1}$ is hyperbolic from $\Delta_{X_{1}}^{m}$. Hence we may assume that $X_{1}$ and $X_{2}$ are both hyperbolic. Next, by definition, there is an exact sequence:
@@
$1\to\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\widetilde{X_{2}}^{m}\bigr)\to\operatorname{Isom}_{\overline{k}/k}\bigl(\widetilde{X_{1}}^{m}/X_{1},\widetilde{X_{2}}^{m}/X_{2}\bigr)\to\operatorname{Isom}_{k}(X_{1},X_{2})\to 1.$
@@
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\widetilde{X_{2}}^{m}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{\overline{k}/k}\bigl(\widetilde{X_{1}}^{m}/X_{1},\widetilde{X_{2}}^{m}/X_{2}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{k}(X_{1},X_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Inn}(\Delta_{X_{2}}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{G_{k}}^{\operatorname{Out},(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
By the definition of $\widetilde{X_{2}}^{m}\to X_{2}$, we have a canonical identification
$\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\widetilde{X_{2}}^{m}\bigr)\,\cong\,\Delta_{X_{2}}^{m}.$
By Theorem 2.9, the group $\Delta_{X_{2}}^{m}$ is center-free. Therefore,
$\ker(\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\widetilde{X_{2}}^{m}\bigr)\twoheadrightarrow\operatorname{Inn}(\Delta_{X_{2}}^{m}))=\mathrm{C}_{\Pi_{X_{2}}^{(m)}}\left(\Delta_{X_{2}}^{m}\right)$
is trivial. Hence the left-hand vertical arrow in the above commutative diagram is bijective. Moreover, the right-hand vertical arrow is surjective by [7, Theorem 18.1], and injective by [16, Lemma 4.9]. (Note that [16, Lemma 4.9] assumed that $k$ is a field finitely generated over $\mathbb{Q}$. However, the proof can be applied to the case where $k$ is a sub-$p$-adic field.) Thus, by the snake lemma, the middle vertical arrow is also bijective. This completes the proof. ∎
@@
In Theorem 2.11, we assumed that $\Sigma=\{p\}$. If we further assume that $m\geq 3$, this assumption can be weakened to $p\in\Sigma$. The author expects that the same statement should hold for such $\Sigma$ even when $m=2$. However, to prove this, we would need to check whether the proof of [7, Theorem 18.1] applies in this setting as well. At the time of writing, the author has not attempted this modification.